A solution of carbonic acid had a PH of 3.53. The pKa's for this acid are 6.37 and 10.25. What was the initial concentration of the solution?
Answer #1

pH of H2CO3 solution is less than pKa1= 6.37 . Hence, major species at pH = 3.53 will be acidic form i e H2CO3.

Consider a reaction, H2CO3 (aq) + H2O (l) \rightleftharpoons H3O + (aq) + HCO 3- (aq)

For above reaction, Ka1 = [H3O +] [HCO 3-] / [H2CO3]

Where, Concentrations of H2CO3 , H3O + and HCO 3- are equilibrium concentrations.

We can calculate equilibrium [H3O +] from pH.

We have , pH = - log [H3O +]

\therefore [H3O +] = 10 - pH = 10 - 3.53 = 2.95 \times 10 -04 M

We have, pKa1 = - log Ka1

\therefore Ka1 = 10 - pKa1 = 10 - 6.37 = 4.26 \times 10 -07

From above values , we can write

4.27 \times 10 -07= (2.95 \times 10 -04 )(2.95 \times 10 -04) / [H2CO3]

[H2CO3] = (2.95 \times 10 -04 )(2.95 \times 10 -04) / 4.26 \times 10 -07 = 0.2038 M

i e Equilibrium [H2CO3] = 0.2038 M = Original concentration of [H2CO3] - concentration of dissociated [H2CO3]

= Original concentration of [H2CO3] - [H3O +]

Original concentration of [H2CO3] = Equilibrium [H2CO3] + [H3O +] = 0.2038 + (2.95 \times 10 -04 ) = 0.204 M

ANSWER : Original concentration of [H2CO3] = 0.204 M